You are dealt one card from a 52 card deck.Find the probability that you are dealt a six or a black card.
Answer by fractalier(6550)

P(six or black) = P(six) + P(black) - P(six)*P(black)
= 4/52 + 13/52 - (4/52)*(13/52) = 16/52 = 4/13